# Towards LLM Unlearning Resilient to Relearning Attacks: A Sharpness-Aware Minimization Perspective and Beyond

### v2

@@ 2 Related Work
Machine unlearning modifies models to remove the influence of undesirable data, originally developed to mitigate post-training privacy risks (Cao & Yang, 2015; Ginart et al., 2019; Ullah et al., 2021; Fan et al., 2024c). While retraining from scratch guarantees exact unlearning, it is computationally prohibitive, leading to research on approximate unlearning methods that balance efficiency and effectiveness (Kurmanji et al., 2024; Fan et al., 2024a; Chen et al., 2023; Zhang et al., 2024c; d). A rapidly growing subfield is LLM unlearning (Jang et al., 2022; Meng et al., 2022; Yao et al., 2023; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024a; Maini et al., 2024; Jia et al., 2024a; Liu et al., 2024c; Fan et al., 2024b; Thaker et al., 2024), which has been shown promise in mitigating the generation of harmful content (Yao et al., 2023; Li et al., 2024; Jia et al., 2024b) and protecting sensitive, copyrighted, or private information (Eldan & Russinovich, 2023; Wu et al., 2023; Jang et al., 2022). Existing LLM unlearning approaches include model-based optimization (Maini et al., 2024; Yao et al., 2023; Jia et al., 2024a; Fan et al., 2024b; Zhang et al., 2024a; Li et al., 2024; Jia et al., 2024a; Wu et al., 2023; Fan et al., 2024b) and input-based strategies (via prompting or in-context learning) to facilitate unlearning without extensive parameter adjustments (Liu et al., 2024a; Thaker et al., 2024; Pawelczyk et al., 2023). Furthermore, recent benchmarking efforts provide valuable frameworks for evaluating the effectiveness of LLM unlearning approaches. These include TOFU (Maini et al., 2024), which focuses on fictitious unlearning using synthetic data, WMDP (Li et al., 2024), which aims to mitigate sociotechnical harms in model generation, and MUSE (Shi et al., 2024), which focuses on erasing copyrighted information from LLMs.
@@
Sharpness-aware minimization (SAM) is a representative smoothness optimization technique that minimizes both the loss value and its sharpness, effectively promoting a flatter loss landscape, originally introduced to improve model generalization (Foret et al., 2021; Andriushchenko & Flammarion, 2022; Liu et al., 2022b; Du et al., 2022; Zhang et al., 2023). SAM has also been applied in traditional adversarial training to defend against input-level adversarial attacks (Wei et al., 2023; Zhang et al., 2024b). Beyond SAM, other smoothness optimization approaches include gradient penalty (GP) and curvature regularization (CR), which impose penalties based on loss gradients or Hessian-gradient products to encourage smoothness (Dauphin et al., 2024; Zhao et al., 2024). Randomized smoothing (RS) improves smoothness by convolving a non-smooth objective function with a Gaussian distribution (Duchi et al., 2012; Cohen et al., 2019; Ji et al., 2024a). Meanwhile, weight averaging (WA) enhances smoothness by averaging model weights across training iterations, leading to a smoother optimization trajectory (Izmailov et al., 2018). Further, activation smoothing has been used to mitigate the security risk of users fine-tuning for LLMs (Huang et al., 2024). These smoothness optimization approaches will serve as a key foundation for enhancing the robustness of LLM unlearning in this work.

### v3

@@ 2 Related Work
Machine unlearning modifies models to remove the influence of undesirable data, originally developed to mitigate post-training privacy risks (Cao & Yang, 2015; Ginart et al., 2019; Ullah et al., 2021; Fan et al., 2024c). While retraining from scratch guarantees exact unlearning, it is computationally prohibitive, leading to research on approximate unlearning methods that balance efficiency and effectiveness (Kurmanji et al., 2024; Fan et al., 2024a; Chen et al., 2023; Zhang et al., 2024c; d). A rapidly growing subfield is LLM unlearning (Jang et al., 2022; Meng et al., 2022; Yao et al., 2024; Eldan & Russinovich, 2023; Jia et al., 2024b; Zhang et al., 2024a; Maini et al., 2024; Jia et al., 2024a; Liu et al., 2024c; Fan et al., 2024b; Thaker et al., 2024), which has been shown promise in mitigating the generation of harmful content (Yao et al., 2024; Li et al., 2024; Jia et al., 2024b) and protecting sensitive, copyrighted, or private information (Eldan & Russinovich, 2023; Wu et al., 2023; Jang et al., 2022). Existing LLM unlearning approaches include model-based optimization (Maini et al., 2024; Yao et al., 2024; Jia et al., 2024a; Fan et al., 2024b; Zhang et al., 2024a; Li et al., 2024; Jia et al., 2024a; Wu et al., 2023; Fan et al., 2024b) and input-based strategies (via prompting or in-context learning) to facilitate unlearning without extensive parameter adjustments (Liu et al., 2024a; Thaker et al., 2024; Pawelczyk et al., 2023). Furthermore, recent benchmarking efforts provide valuable frameworks for evaluating the effectiveness of LLM unlearning approaches. These include TOFU (Maini et al., 2024), which focuses on fictitious unlearning using synthetic data, WMDP (Li et al., 2024), which aims to mitigate sociotechnical harms in model generation, and MUSE (Shi et al., 2024), which focuses on erasing copyrighted information from LLMs.
@@
Sharpness-aware minimization (SAM) is a representative smoothness optimization technique that minimizes both the loss value and its sharpness, effectively promoting a flatter loss landscape, originally introduced to improve model generalization (Foret et al., 2021; Andriushchenko & Flammarion, 2022; Liu et al., 2022b; Du et al., 2022; Zhang et al., 2023; Jiang et al., 2023). SAM has also been applied in traditional adversarial training to defend against input-level adversarial attacks (Wei et al., 2023; Zhang et al., 2024b). Beyond SAM, other smoothness optimization approaches include gradient penalty (GP) and curvature regularization (CR), which impose penalties based on loss gradients or Hessian-gradient products to encourage smoothness (Dauphin et al., 2024; Zhao et al., 2024). Randomized smoothing (RS) improves smoothness by convolving a non-smooth objective function with a Gaussian distribution (Duchi et al., 2012; Cohen et al., 2019; Ji et al., 2024a). Meanwhile, weight averaging (WA) enhances smoothness by averaging model weights across training iterations, leading to a smoother optimization trajectory (Izmailov et al., 2018). Further, activation smoothing has been used to mitigate the security risk of users fine-tuning for LLMs (Huang et al., 2024). These smoothness optimization approaches will serve as a key foundation for enhancing the robustness of LLM unlearning in this work.